# Quantitative 2D propagation of smallness and control for 1D heat equations with power growth potentials

### v1

@@ 2.3. Reduction of potentials
(2.25) $\partial_{t}U-\partial_{x}^{2}U+V(x)U+\theta U=0,\quad U\lvert_{t=0}=u_{0}\in L^{2}(\mathbb{R}).$
@@ 3. Exact propagation of smallness
for some positive numerical constants $d_{1}$ and $d_{2}$.
@@ 3.1. Preliminaries
Given any $\varepsilon>0$ and any sequence of balls $\left\{B_{r_{i}}(z_{i})\right\}_{i\in\mathcal{I}}$ covering $E$ such that $\sum_{i\in\mathcal{I}}r^{\kappa\delta}\leq\mathcal{H}_{\kappa\delta}(E)+\varepsilon$, then $\left\{G\left(B_{r_{i}}(z_{i})\right)\right\}_{i\in\mathcal{I}}$ is a covering of $G(E)$. By the Hölder continuity of $G$, we obtain $\left\{B_{r_{i}^{\prime}}(G(z_{i}))\right\}$ the covering of $G(E)$ where $r_{i}^{\prime}:=Cr^{\kappa}$ for all $i\in\mathcal{I}$. Then we obtain
(3.17) $\mathcal{H}_{\delta}\left(G(E)\right)\leq\sum_{i\in\mathcal{I}}r_{i}^{\prime\delta}=\sum_{i\in\mathcal{I}}C^{\delta}r^{\kappa\delta}\leq C^{\delta}\mathcal{H}_{\kappa\delta}(E)+\varepsilon.$
@@
Let $u$ be harmonic in an open connected set $\Omega$ in $\mathbb{R}^{2}$ and let Ωsuperscriptnormal-$\Omega^{\prime}$ be any compact subset of $\Omega$. Then for any multi-index $\alpha$ we have
@@
Let $\omega\subset B_{1}\cap\ell_{0}$ satisfy $|\omega|>0$ for some line $\ell_{0}$ in $\mathbb{R}^{2}$. Then for some $z_{0}\in\omega$, there exist some constants $\alpha=\alpha(\Lambda,|\omega|)\in(0,1)$ and $C(\Lambda,|\omega|)>0$, such that for any real-valued $H^{2}_{\mathrm{loc}}$ solution $\phi$ of (3.1) with its $A$-harmonic conjugate satisfying $\psi_{z_{0}}(z_{0})=0$, we have
@@
We choose
@@
(3.34) $|G_{a}(z)|=|G_{a}(z)-G_{a}(a)|\geq\left(\frac{1}{32}\right)^{2(1+K)K}|z-a|\geq\left(\frac{1}{32}\right)^{2(1+K)K}\frac{|\omega|}{8}.$
@@
(3.35) $6r_{2}=\left(\frac{1}{32}\right)^{2(1+K)K}\frac{|\omega|}{8}.$
@@
(3.37) $|\omega\cap B_{r_{3}}(z_{0})|\geq\left(\frac{1}{32}\right)^{2(1+K)^{3}}|\omega|^{K+1}$
@@
(3.40) $r_{3}=\left(\frac{1}{32}\right)^{2(1+K)K}r_{2}^{K}=\left(\frac{1}{48}\right)^{K}\left(\frac{1}{32}\right)^{2(1+K)^{2}K}|\omega|^{K}.$
@@
(3.41) $|\omega\cap\ell_{n_{0}}|\geq\frac{|\omega|}{2\lceil\frac{2r_{1}}{r_{3}}\rceil}.$
Choose one $z_{0}\in\ell_{n_{0}}$ and then we have
@@
(3.43) $2\lceil\frac{2r_{1}}{r_{3}}\rceil\leq 4\times 48^{K}\times 32^{2(1+K)^{2}K}\,\frac{2-\frac{|\omega|}{4}}{|\omega|^{K}}\leq{32^{2(1+K)^{3}}}|\omega|^{-K}.$
Combining this with (3.42) we obtain the desired (3.37), proving the claim. ∎
@@
Let $a_{1},a_{2},\cdots,a_{N}$ be the zeros of $F_{z_{0}}$ in $B_{2r_{2}}$ (repeated according to their respective multiplicity). Let $a_{0}\in\partial B_{r_{2}}$ be such that $|F_{z_{0}}(a_{0})|$ is the maximum of $|F_{z_{0}}(z)|$ on $B_{r_{2}}$. We consider the following complex polynomial $P_{z_{0}}(z)$ and holomorphic nonvanishing function $h_{z_{0}}(z)$ on $B_{2r_{2}}$,
@@
(3.48) $\sup_{B_{r_{2}}}\left(\sup_{B_{5r_{2}}}\log|h_{z_{0}}|-\log|h_{z_{0}}(z)|\right)\leq 9\inf_{B_{r_{2}}}\left(\sup_{B_{5r_{2}}}\log|h_{z_{0}}|-\log|h_{z_{0}}(z)|\right).$
It follows that
@@
On the other hand, using Lemma 3.9, we find
@@
(3.54) $\sup_{B_{r_{2}}}|P_{z_{0}}|$ $\leq\left(\frac{3e\times 32^{2(1+K)^{3}+2(1+K)}}{|\omega|^{K+1}}\right)^{KN}\sup_{G_{z_{0}}(\widetilde{\omega})}|P_{z_{0}}|$
$\leq\left(\frac{32^{2(1+K)^{4}}}{|\omega|^{K+1}}\right)^{KN}\sup_{G_{z_{0}}(\widetilde{\omega})}|P_{z_{0}}|.$
@@
$\leq$ $\left(\frac{32^{2(1+K)^{4}K}}{|\omega|^{(K+1)K}}\right)^{N}\left(\sup_{B_{5r_{2}}}|F_{z_{0}}|^{9}\right)\left(\inf_{B_{r_{2}}}|h_{z_{0}}|\right)\left(\sup_{G_{z_{0}}(\widetilde{\omega})}|P_{z_{0}}|\right)$
$\leq$ $\left(\frac{32^{2(1+K)^{4}K}}{|\omega|^{(K+1)K}}\right)^{N}\left(\sup_{B_{5r_{2}}}|F_{z_{0}}|^{9}\right)\left(\sup_{G_{z_{0}}(\widetilde{\omega})}|F_{z_{0}}|\right)$
@@
(3.59) $\alpha_{1}:=$ $\frac{1}{10+\frac{\log C_{0}}{\log(7/6)}}=\frac{\log(7/6)}{10\log(7/6)+2(1+K)^{4}K\log 32-(K+1)K\log|\omega|}$
$\asymp$ $\frac{1}{K^{2}(K^{3}+\bigl{|}\log|\omega|\bigr{|})}.$
@@
(3.61) $\alpha_{2}=\frac{\log\frac{2+r}{2r}}{\log\frac{2+r}{2r_{2}}}={\log\frac{4-\left(\frac{1}{32}\right)^{2(1+K)K}}{4-2\left(\frac{1}{32}\right)^{2(1+K)K}}}\bigg{/}{\log\frac{4-\left(\frac{1}{32}\right)^{2(1+K)K}}{\frac{|\omega|}{24}\left(\frac{1}{32}\right)^{2(1+K)K}}}\asymp\frac{32^{-2(1+K)K}}{K^{2}+\bigl{|}\log|\omega|\bigr{|}}.$
@@
Combining the above inequality (3.62) with $\widetilde{\omega}\subset\omega$, we obtain
@@
Let $z_{0}\in\omega$ be the point chosen in Proposition 3.11 and $\psi_{z_{0}}$ be the $A$-harmonic conjuegate of $\Psi$ satisfying $\psi_{z_{0}}(z_{0})=0$. By $A\nabla\Psi_{z_{0}}\cdot\textbf{e}_{0}=0$ and the definition of $\psi_{z_{0}}$, we obtain
@@ 4.2. Propagation of smallness: nondivergence form
Remember from Subsection 2.1 that $\{x_{n}\}_{n\in\mathbb{N}}$ is defined by (2.1), $I_{n}=[x_{n},x_{n+1}]$ for all $n\in\mathbb{N}_{+}$, $I_{0}=[-x_{0},x_{0}],\,x_{0}=1$ and $I_{-n}=-I_{n}$ for all $n\in\mathbb{N}$. We define
(4.11) $I_{2,n}:=\left[x_{n}-|I_{n}|,x_{n+1}+|I_{n}|\right],\quad\forall n\in\mathbb{N}$
@@
(4.12) $I_{3,n}:=\left[x_{n}-2|I_{n}|,x_{n+1}+2|I_{n}|\right],\quad\forall n\in\mathbb{N}.$
We also denote by $I_{2,n}:=-I_{2,|n|}$ and $I_{3,n}:=-I_{3,|n|}$ for all $n\in-\mathbb{N}_{+}$. Define
[…]
We also define
[…]
@@
The following lemma can be found in [SSY23, Lemma 3.1]:
@@
Let $I=[a,b]$ be a finite interval and let $V\in C(I)\cap L^{\infty}(I)$ with $V>0$. Then there exists a $\mathcal{C}^{2}$ positive solution $\varphi$ of (4.15) on the interval $I$ such that
(4.16) $1\leq\varphi\leq e^{(b-a)^{2}\|V\|_{L^{\infty}(I)}},\quad\forall x\in I.$
Remark 4.4.
Notice that in [SSY23, Lemma 3.1] the condition posed on $V$ is $V\geq 1$, which makes no difference here. Lemma 4.3 explains why we assume the continuity for the potential $V$: to ensure the existence and regularity of the solution for (4.15).
@@
(4.17) $-\Delta\Phi(z)+V(x)\Phi(z)=0\,\text{ with }\,\partial_{y}\Phi\lvert_{y=0}.$
@@
(4.18) $\begin{cases}\partial_{y}\Phi|_{y=0}&\Longrightarrow\partial_{y}\left.\left(\frac{\Phi(z)}{\varphi(x)}\right)\right|_{y=0}=0,\\ -\Delta\Phi(z)+V(x)\Phi(z)=0&\Longrightarrow\nabla\cdot\left(\varphi^{2}\nabla\bigl{(}\frac{\Phi(z)}{\varphi(x)}\bigr{)}\right)=0.\end{cases}$
@@
Let $C_{0}>0$ be a positive constant. Then for any measurable set $\omega\subset I=[0,1]$ with $|\omega|\in(0,\frac{1}{2})$, any potential $V\in\mathcal{C}(\mathbb{R})$ with $0<V(x)\leq C_{0}$ for any $x\in I_{4}=[-4\sqrt{2},4\sqrt{2}]$, and any real-valued $H^{2}_{\mathrm{loc}}$ solution $\Phi$ of (4.17) in $\mathbb{R}^{2}$, we have
@@
(4.21) $\alpha\geq\frac{\exp\bigl{(}-d\exp(dC_{0})\bigr{)}}{\bigl{|}\log|\omega|\bigr{|}^{2}},.$
@@
where $\alpha$ and $C$ satisfy (3.4) and (3.5) where $\Lambda=e^{d_{3}C_{0}}$, and $d_{1},d_{2},d_{3}$ are numerical positive constants. The numerical constants may be different from line to line. Indeed, we consider the positive $\mathcal{C}^{2}$ solution $\varphi$ constructed in Lemma 4.3 for the second-order ODE (4.15) on $I_{4}$. Then from the reduction (4.18), we can apply Theorem 3.1 to $(\Phi/\varphi)$ with $\ell_{0}=\left\{(x,y)\in\mathbb{R}^{2}:y=0\right\}$. According to the scaling method, we first replace $B_{1}$ and $B_{2}$ in Theorem 3.1 with $B_{1}^{\prime}:=B_{\frac{\sqrt{2}}{2}}\left(\frac{1}{2}\right)$ and $B_{2}^{\prime}:=B_{\sqrt{2}}\left(\frac{1}{2}\right)$, then we have
@@
where $\alpha$ and $C$ satisfy (3.4) and (3.5) with $\Lambda=e^{d_{3}C_{0}}$ by Lemma 4.3, $d_{1},d_{2},d_{3}>0$ are numerical positive constants and $d_{1},d_{2}$ may be possibly different from those in (3.4) and (3.5).
@@
(4.25) $\sup_{D_{1}}|\Phi|\leq Ce^{C_{0}}\left(\sup_{\omega}|\Phi|^{\alpha}\right)\left(\sup_{D_{2}}|\Phi|^{1-\alpha}\right),$
Replacing $Ce^{C_{0}}$ with a new $C$ and $d_{1},d_{2},d_{3}$ with a new $d$ large enough, we complete the proof of (4.22) with the constants given by (3.4) and (3.5).

### v2

@@ 2.3. Reduction of potentials
(2.26) $\partial_{t}U-\partial_{x}^{2}U+V(x)U+\theta U={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}e^{\theta t}h(t,x)\mathbbm{1}_{\Omega}},\quad U\lvert_{t=0}=u_{0}\in L^{2}(\mathbb{R}).$
@@ 3. Exact propagation of smallness
for some positive constants $d_{1}$ and $d_{2}$ which depend on the line $\ell_{0}$.
@@ 3.1. Preliminaries
Given any $\varepsilon>0$ and any sequence of balls $\left\{B_{r_{i}}(z_{i})\right\}_{i\in\mathcal{I}}$ covering $E$ such that $\sum_{i\in\mathcal{I}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r_{i}}^{\kappa\delta}\leq\mathcal{H}_{\kappa\delta}(E)+\varepsilon$, then $\left\{G\left(B_{r_{i}}(z_{i})\right)\right\}_{i\in\mathcal{I}}$ is a covering of $G(E)$. By the Hölder continuity of $G$, we obtain $\left\{B_{r_{i}^{\prime}}(G(z_{i}))\right\}$ the covering of $G(E)$ where $r_{i}^{\prime}:={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Cr_{i}^{\kappa}}$ for all $i\in\mathcal{I}$. Then we obtain
(3.17) $\mathcal{H}_{\delta}\left(G(E)\right)\leq\sum_{i\in\mathcal{I}}r_{i}^{\prime\delta}=\sum_{i\in\mathcal{I}}C^{\delta}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r_{i}}^{\kappa\delta}\leq C^{\delta}\mathcal{H}_{\kappa\delta}(E)+\varepsilon.$
@@
Let $u$ be harmonic in an open connected set $\Omega$ in $\mathbb{R}^{2}$ and let $\Omega^{\prime}$ be any compact subset of $\Omega$. Then for any multi-index $\alpha$ we have
@@
Let $\omega\subset B_{1}\cap\ell_{0}$ satisfy $|\omega|>0$ for the line $\ell_{0}=\{(x,y)\in\mathbb{R}^{2}:y=0\}$. Then for some $z_{0}\in\omega$, there exist some constants $\alpha=\alpha(\Lambda,|\omega|)\in(0,1)$ and $C(\Lambda,|\omega|)>0$, such that for any real-valued $H^{2}_{\mathrm{loc}}$ solution $\phi$ of (3.1) with its $A$-harmonic conjugate satisfying $\psi_{z_{0}}(z_{0})=0$, we have
@@
The strategy of the proof is to decompose the function $f_{z_{0}}=\phi+i\psi_{z_{0}}$ into $f_{z_{0}}=F_{z_{0}}\circ G_{z_{0}}$ for some $z_{0}\in B_{1}$, and reduce it to the holomorphic case since $F$ is a holomorphic function, that is, reduce to the inequality
(3.32) $\sup_{G_{z_{0}}(B_{1})}|F_{z_{0}}|\leq C\left(\sup_{G_{z_{0}}(\omega)}|F_{z_{0}}|^{\alpha}\right)\left(\sup_{B_{2}}|\phi|^{1-\alpha}\right)$
for some $C>0$ and $\alpha\in(0,1)$. To achieve this, we need to make sure that the quasiconformal mapping $G_{z_{0}}$ will not change the “shape” of the domain too much, which is controlled by Mori’s Theorem.
First, we need to choose a ball $B_{r_{1}}\subset B_{1}$ with $r_{1}<1$ so that the “shape” of $G_{a}(B_{1})$ for any $a\in B_{r_{1}}$ can be uniformly controlled and $\omega$ restricted to this ball has positive $1$-dimensional Lebesgue measure. Otherwise, if we treat any $a\in B_{1}$ directly, we cannot have a uniform estimate for the “shape” of $G_{a}(B_{1})$ through Mori’s Theorem. This kind of $r_{1}$ can be chosen as $r_{1}=1-a|\omega|$ for any $a\in(0,\frac{1}{2})$. For simplicity, we choose
@@
(3.35) $|G_{a}(z)|=|G_{a}(z)-G_{a}(a)|\geq\left(\frac{1}{32}\right)^{2(1+K)K}|z-a|^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}}\geq\left(\frac{1}{32}\right)^{2(1+K)K}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\frac{|\omega|}{8}\right)^{K}}.$
@@
(3.36) $6r_{2}=\left(\frac{1}{32}\right)^{2(1+K)K}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\frac{|\omega|}{8}\right)^{K}}.$
@@
(3.38) $|\omega\cap B_{r_{3}}(z_{0})|\geq\left(\frac{1}{32}\right)^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}(1+K)^{3}}|\omega|^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K^{2}}+1}$
@@
(3.41) $r_{3}=\left(\frac{1}{32}\right)^{2(1+K)K}r_{2}^{K}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\frac{1}{6^{K}8^{K^{2}}}\right)\left(\frac{1}{32}\right)^{2(1+K)^{2}K}|\omega|^{K^{2}}.}$
@@
(3.42) $|\omega\cap\ell_{n_{0}}|\geq\frac{|\omega|}{2\lceil\frac{2r_{1}}{r_{3}}\rceil},$
where we used the fact $|\omega\cap B_{r_{1}^{\prime}}|\geq|\omega|-2(1-r_{1}^{\prime})=\frac{|\omega|}{2}$. Choose one $z_{0}\in\ell_{n_{0}}$ and then we have
@@
(3.44) $2\lceil\frac{2r_{1}}{r_{3}}\rceil\leq 4\times 48^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K^{2}}}\times 32^{2(1+K)^{2}K}\,\frac{2-\frac{|\omega|}{4}}{|\omega|^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K^{2}}}\leq{32^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}(1+K)^{3}}}|\omega|^{-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K^{2}}}$
where we used the facts $6^{K}8^{K^{2}}<48^{K^{2}}$, $2-\frac{|\omega|}{4}\leq 2$ and then $8\times 48^{K^{2}}\times 32^{2(1+K)^{2}K}\leq 32^{4(1+K)^{3}}$. Combining this with (3.43) we obtain the desired (3.38), proving the claim. ∎
@@
Let $a_{1},a_{2},\cdots,a_{N}$ be the zeros of $F_{z_{0}}$ in $B_{2r_{2}}$ (repeated according to their respective multiplicity). Let $a_{0}\in\partial B_{r_{2}}$ be such that $|F_{z_{0}}(a_{0})|$ is the maximum of $|F_{z_{0}}(z)|$ on $B_{r_{2}}$. We consider the following complex polynomial $P_{z_{0}}(z)$ and holomorphic function $h_{z_{0}}(z)$ which is nonvanishing on $B_{2r_{2}}$,
@@
(3.49) $\sup_{B_{r_{2}}}\left(\sup_{B_{5r_{2}}}\log|h_{z_{0}}|-\log|h_{z_{0}}(z)|\right)\leq 9\inf_{B_{r_{2}}}\left(\sup_{B_{5r_{2}}}\log|h_{z_{0}}|-\log|h_{z_{0}}(z)|\right),$
where the constant comes from $\left(\frac{R+r}{R-r}\right)^{2}=9$ by choosing $R=2r_{2}$ and $r=r_{2}$. It follows that
@@
On the other hand, using Lemma 3.9 and the inclusion condition (3.39), we find
@@
(3.55) $\sup_{B_{r_{2}}}|P_{z_{0}}|$ $\leq\left(\frac{3e\times 32^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}(1+K)^{3}+2(1+K)}}{|\omega|^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K^{2}}+1}}\right)^{KN}\sup_{G_{z_{0}}(\widetilde{\omega})}|P_{z_{0}}|$
$\leq\left(\frac{32^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}(1+K)^{4}}}{|\omega|^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K^{2}}+1}}\right)^{KN}\sup_{G_{z_{0}}(\widetilde{\omega})}|P_{z_{0}}|.$
@@
$\leq$ $\left(\frac{32^{4(1+K)^{4}K}}{|\omega|^{(K^{2}+1)K}}\right)^{N}\left(\sup_{B_{5r_{2}}}|F_{z_{0}}|^{9}\right)\left(\inf_{B_{r_{2}}}|h_{z_{0}}|\right)\left(\sup_{G_{z_{0}}(\widetilde{\omega})}|P_{z_{0}}|\right)$
$\leq$ $\left(\frac{32^{4(1+K)^{4}K}}{|\omega|^{(K^{2}+1)K}}\right)^{N}\left(\sup_{B_{5r_{2}}}|F_{z_{0}}|^{9}\right)\left(\sup_{G_{z_{0}}(\widetilde{\omega})}|F_{z_{0}}|\right)$
@@
(3.60) $\alpha_{1}:=$ $\frac{1}{10+\frac{\log C_{0}}{\log(7/6)}}=\frac{\log(7/6)}{10\log(7/6)+4(1+K)^{4}K\log 32-(K^{2}+1)K\log|\omega|}$
$\asymp$ $\frac{1}{K^{3}(K^{2}+\bigl{|}\log|\omega|\bigr{|})}.$
@@
(3.62) $\alpha_{2}=\frac{\log\frac{2+r}{2r}}{\log\frac{2+r}{2r_{2}}}={\log\frac{4-\left(\frac{1}{32}\right)^{2(1+K)K}}{4-2\left(\frac{1}{32}\right)^{2(1+K)K}}}\bigg{/}{\log\frac{4-\left(\frac{1}{32}\right)^{2(1+K)K}}{\frac{1}{3}\left(\frac{|\omega|}{8}\right)^{K}\left(\frac{1}{32}\right)^{2(1+K)K}}}\asymp\frac{32^{-2(1+K)K}}{K^{2}+K\bigl{|}\log|\omega|\bigr{|}}.$
@@
Combining the above inequality with (3.63) and noticing that $\widetilde{\omega}\subset\omega$, we obtain
@@
Without loss of generality, we choose the line to be the standard case $\ell_{0}=\{(x,y)\in\mathbb{R}^{2}:y=0\}$. Indeed, for arbitrary line, we can choose a smaller ball $B_{r_{0}}$ which has the center located in the line and change the coordinate system to make it be horizontal. By using Hadamard’s three-circle theorem for elliptic equation with the smallest ball being $B_{r_{0}}$, the left of the proof is the same as the standard case. Let $z_{0}\in\omega$ be the point chosen in Proposition 3.11 and $\psi_{z_{0}}$ be the $A$-harmonic conjuegate of $\Psi$ satisfying $\psi_{z_{0}}(z_{0})=0$. By $A\nabla\Psi_{z_{0}}\cdot\textbf{e}_{0}=0$ and the definition of $\psi_{z_{0}}$, we obtain
@@ 4.2. Propagation of smallness: nondivergence form
Remember from Subsection 2.1 that $\{x_{n}\}_{n\in\mathbb{N}}$ is defined by (2.1), $I_{n}=[x_{n},x_{n+1}]$ for all $n\in\mathbb{N}_{+}$ and $I_{0}=[-x_{0},x_{0}],\,x_{0}=1$, $I_{-n}=-I_{n}$ for all $n\in\mathbb{N}$. We define
(4.11) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{1,n}:=I_{n},\,I_{2,n}:=\left[x_{n}-|I_{n}|,x_{n+1}+|I_{n}|\right],\,I_{3,n}:=\left[x_{n}-2|I_{n}|,x_{n+1}+2|I_{n}|\right],\quad\forall n\in\mathbb{N}.}$
We also denote by $I_{j,n}:=-I_{j,|n|}$ for all $n\in-\mathbb{N}_{+}$ and $j=1,2,3$. Define
(4.12) $D_{j,n}:=I_{j,n}\times\left[-\frac{(2j-1)|I_{n}|}{2},\frac{(2j-1)|I_{n}|}{2}\right],\quad\forall n\in\mathbb{N},j=1,2,3,$
@@
(4.13) $D_{j}:=[1-j,j]\times\left[-\frac{(2j-1)}{2},\frac{(2j-1)}{2}\right],\quad\forall n\in\mathbb{N},j=1,2,3.$
@@
We have the following lemma which is a particular case of the elliptic theory and can be found in [LM24, Lemma 3.1]:
@@
Let $I=[a,b]$ be a finite interval and let $V\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{\infty}}(I)$ with $V\geq 0$. Then there exists a positive solution $\varphi\in H^{2}([a,b])$ of (4.14) on the interval $I$ such that
(4.15) $1\leq\varphi\leq e^{(b-a)\|V\|_{L^{\infty}(I)}^{\frac{1}{2}}},\quad\forall x\in I.$
Proof.
We first solve a second-order ODE problem with a boundary condition, by [GT77, Theorem 9.15], there exists $\varphi\in H^{2}([a,b])$ such that
(4.16) $-\varphi^{\prime\prime}(x)+V(x)\varphi(x)=0,\quad x\in[a,b],\,\varphi(a)=\varphi(b)=e^{(b-a)\|V\|_{L^{\infty}(I)}^{\frac{1}{2}}}.$
Define
(4.17) $\phi_{-}(x)=e^{(b-x)\|V\|_{L^{\infty}(I)}^{\frac{1}{2}}}\text{ and }\phi_{+}(x)=e^{(b-a)\|V\|_{L^{\infty}(I)}^{\frac{1}{2}}},\quad\forall x\in[a,b].$
Then we have
(4.18) $-\phi_{-}^{\prime\prime}+V\phi_{-}\leq 0,\quad\forall x\in(a,b),\quad\phi_{-}(a),\phi_{-}(b)\leq e^{(b-a)\|V\|_{L^{\infty}(I)}^{\frac{1}{2}}},$
and
(4.19) $-\phi_{+}^{\prime\prime}+V\phi_{+}\geq 0,\quad\forall x\in(a,b),\quad\phi_{+}(a),\phi_{+}(b)\geq e^{(b-a)\|V\|_{L^{\infty}(I)}^{\frac{1}{2}}}.$
So by the maximal principle stated in [GT77, Theorem 8.1], we have
(4.20) $\phi_{-}\leq\phi(x)\leq\phi_{+},\quad\forall x\in(a,b).$
This implies (4.15) and finishes the proof. ∎
@@
(4.21) $-\Delta\Phi(z)+V(x)\Phi(z)=0\,\text{ with }\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\partial_{y}\Phi\lvert_{y=0}=0}.$
@@
(4.22) $\begin{cases}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\partial_{y}\Phi|_{y=0}=0}&\Longrightarrow\partial_{y}\left.\left(\frac{\Phi(z)}{\varphi(x)}\right)\right|_{y=0}=0,\\ -\Delta\Phi(z)+V(x)\Phi(z)=0&\Longrightarrow\nabla\cdot\left(\varphi^{2}\nabla\bigl{(}\frac{\Phi(z)}{\varphi(x)}\bigr{)}\right)=0.\end{cases}$
@@
Let $C_{0}>0$ be a positive constant. Then for any measurable set $\omega\subset I=[0,1]$ with $|\omega|\in(0,\frac{1}{2})$, any potential $V\in L^{\infty}(\mathbb{R})$ with $0<V(x)\leq C_{0}$ for any $x\in I_{4}=[-4\sqrt{2},4\sqrt{2}]$, and any real-valued $H^{2}_{\mathrm{loc}}$ solution $\Phi$ of (4.21) in $\mathbb{R}^{2}$, we have
@@
(4.25) $\alpha\geq\frac{\exp\bigl{(}-d\exp(dC_{0}^{\frac{1}{2}})\bigr{)}}{\bigl{|}\log|\omega|\bigr{|}^{2}},.$
@@
where $\alpha$ and $C$ satisfy (3.4) and (3.5) where $\Lambda=e^{d_{3}C_{0}^{\frac{1}{2}}}$, and $d_{1},d_{2},d_{3}$ are numerical positive constants. The numerical constants may be different from line to line. Indeed, we consider the positive $W^{1,\infty}$ solution $\varphi$ constructed in Lemma 4.3 for the second-order ODE (4.14) on $I_{4}$. Then from the reduction (4.22), we can apply Theorem 3.1 to $(\Phi/\varphi)$ with $\ell_{0}=\left\{(x,y)\in\mathbb{R}^{2}:y=0\right\}$. According to the scaling method, we first replace $B_{1}$ and $B_{2}$ in Theorem 3.1 with $B_{1}^{\prime}:=B_{\frac{\sqrt{2}}{2}}\left(\frac{1}{2}\right)$ and $B_{2}^{\prime}:=B_{\sqrt{2}}\left(\frac{1}{2}\right)$, then we have
@@
where $\alpha$ and $C$ satisfy (3.4) and (3.5) with $\Lambda=e^{d_{3}C_{0}^{\frac{1}{2}}}$ by Lemma 4.3, $d_{1},d_{2},d_{3}>0$ are numerical positive constants and $d_{1},d_{2}$ may be possibly different from those in (3.4) and (3.5).
@@
(4.29) $\sup_{D_{1}}|\Phi|\leq Ce^{C_{0}^{\frac{1}{2}}}\left(\sup_{\omega}|\Phi|^{\alpha}\right)\left(\sup_{D_{2}}|\Phi|^{1-\alpha}\right),$
where $\alpha$ and $C$ satisfy (3.4) and (3.5) with $\Lambda=e^{d_{3}C_{0}^{\frac{1}{2}}}$ and $d_{1},d_{2},d_{3}>0$. By choosing new $d_{1},d_{2}$ appropriately, we complete the proof of (4.26) with the constants given by the forms which are the same as those in (3.4) and (3.5).